A block with an initial velocity of 8.9 m/s slides across a rough surface until it eventually comes to rest. If the coefficient of friction is equal to 0.83, then how far will the block travel?
1
Expert's answer
2020-02-07T10:54:48-0500

We can find the distance that the block travel until it eventually comes to rest from the kinematic equation:


"v_f^2 - v_i^2 = 2as,"

here, "v_i = 8.9 \dfrac{m}{s}" is the initial velocity of the block, "v_f = 0 \dfrac{m}{s}" is the final velocity of the block, "a" is the acceleration of the block and "s" is the distance that the block travel until it eventually comes to rest.

From this equation we can find the distance that the block travel until it eventually comes to rest:


"s = \dfrac{-v_i^2}{2a}."


We can find the acceleration of the block from the Newton's Second Law of Motion:


"\sum F_x = ma_x""-F_{fr} = ma,""a = \dfrac{-F_{fr}}{m} = \dfrac{- \mu N}{m} = \dfrac{- \mu mg}{m} = - \mu g."

Finally, substituting expression for "a" into the formula for distance, we get:


"s = \dfrac{-v_i^2}{2(- \mu g)} = \dfrac{(8.9 \dfrac{m}{s})^2}{2 \cdot 0.83 \cdot 9.8 \dfrac{m}{s^2}} = 4.87m."

Answer:

"s = 4.87m."